6. (10 points) Find the general solution to the DE using the method of Variation of Parameters. 2y - 4y - 22y +24y = 2e
Answer #1

24 - uy! - 23y +24y = gete y-24-lytrzy =eur fist solve foro Homogenous solutow y-24 -llyltlay=0 03-2D2-1D +/2-0 (0-1)